OpenStudy (anonymous):

How do I solve this equation: (2x)^2=x^2+36^2?

OpenStudy (anonymous):

Meh, I just did the whole thing, but server crashed.. :P Oh well, I'll do it again. (2x)^2 = 4x^2 So you got 4x^2 = x^2+36^2 By subtracting x^2 on both sides you get: 3x^2 = 36^2 Take your calculator, and do 36^2, you'll get 1296 You now have: 3x^2 = 1296 Divide both sides by 3 gives: x^2 = 432 Which gives \[x = +- \sqrt{}\]

OpenStudy (anonymous):

Ofcourse the server bugs again... Continuing where it failed: Which gives: x = +- squareroot of 432